Anna is 30 years older than Corey. In 4 years, Anna will be double Corey's age. How old is each now?
Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52852)

.
A = C + 30         (1)
A + 4 = 2(C+4)     (2)


To solve the system, substitute eq(1) into eq(2).


(C+30) + 4 = 2(C+4)

C + 34 = 2C + 8

34 - 8 = 2C - C

C = 26.


Then from eq(1)  A = 26 + 30 = 56.


Answer.  Anna is 56 years old;  Corey is 26.



Answer by greenestamps(13203)



Anna is 30 years older than Corey; in 4 years she will still be 30 years older.

At that time, she will be twice Corey's age. Since the difference in their ages is still 30 years, Anna will be 60 and Corey will be 30 4 years from now.

So now Anna is 60-4-56 and Corey is 30-4 = 26.